OpenStudy (anonymous):

How would I find the shortest distance between two lines if the lines are in vector form? Do I first put it in parametric form?

OpenStudy (amistre64):

cross them to define a normal and the planes between them is one thought then all you have to do is find the distance between the planes

OpenStudy (amistre64):

spose we have the points (x1, y1, z1) on line 1 and (x2, y2, z2) on line 2 and the direction vectors cross to get (a,b,c) the planes are just: a(x-x1) + b(y-y1) + c(z-z1) = 0 and a(x-x2) + b(y-y2) + c(z-z2) = 0 and the line that stabs between them is some x = x1 + at y = y1 + bt z = z1 + ct insert them into each plane to define 2 points that you can then find the distance between. its not the formula i know but i can never remember the formula

OpenStudy (amistre64):

another idea is after you have the planes .... and a line between them 

OpenStudy (amistre64):

|n| |L| cos(a) = n.L cos(a) = n.L/|n| |L| |L| cos(a) = n.L/|n|

OpenStudy (amistre64):

id go that route, seems simpler lol

OpenStudy (anonymous):

ok thanks!